You bought a stock three months ago for $48.57 per share. The stock paid no dividends. The current share price is $53.09. What is the APR of your investment? The EAR?
Answer #1
According to the given information,
Stock price = $48.57
Current stock price = $53.09
This stock paid no dividend, so the retrun was
R = ($53.09 - $48.57) / 48.57
= 0.093 or 9.3%
For three months, the return was 9.3%
Annual percentage yield (APR) = 4 (9.3%) {Since 3months return is 9.3%, 12months return is calculated as 4 times of 3months return}
= 37.2%
Therefore, the value of APR = 37.2%
Effective Annual Return = [1 + (Annual Rate/ N)] ^N - 1
= [1 + (37.2% / 4)] ^4 - 1
= [ 1 + 0.093] ^4 -1
= [ 1.093]^4 - 1
= 1.4272 - 1
= 0.4272 or 42.72%
Answer #2
According to the given information,
Stock price = $48.57
Current stock price = $53.09
This stock paid no dividend, so the retrun was
                                 R = ($53.09 - $48.57) / 48.57
                                     = 0.093 or 9.3%
For three months, the return was 9.3%
Annual percentage yield (APR) = 4 (9.3%) {Since 3months return is 9.3%, 12months return is calculated as 4 times of 3months return}
                                               = 37.2%
Therefore, the value of APR = 37.2%
Effective Annual Return = [1 + (Annual Rate/ N )] ^N - 1
                                    = [1 + (37.2% / 4)] ^4 - 1
                                    = [ 1 + 0.093] ^4 -1
                                   = [ 1.093]^4 - 1
                                   = 1.4272 - 1
                                   = 0.4272 or 42.72%
Answer #3
APR of Investment = Market Price - Cost Price
= 53.09 - 48.57
= $4.52
In % =4.52/48.57*100
= 9.31%
EAR of Investment = [ 1+ (i/n)]^n - 1
=(1+(0.0931/3))^3-1
= 9.6%
i = APR %
n = Number of items compounded a year.
Thank you....